# Rational Tuning of LLM Cascades via Probabilistic Modeling

### v1

@@ 1 Introduction
As LLMs become workhorses of the modern computing stack, systems of LLMs have received significant attention (Zaharia et al., (2024), Chen et al., 2024b ). These approaches make it possible to adapt computational spending to the performance requirements at the query or task level (Kag et al., (2023), Chen et al., (2023)), yielding significant gains in operational efficiency. These gains are achievable even when accessing LLMs entirely via black-box API calls, by switching between models of different capabilities.
However, moving from single LLMs to LLM systems introduces significant additional complexity. To find the system’s optimal operating point, it is important to understand not just the performance of individual models but also the interactions between their error rates. For example, in a simple two-model LLM cascade in which a small model delegates difficult queries to a large model, the large model’s error rate increases conditional on receiving a query, since the small model’s confidence gating induces an adverse selection (Zellinger and Thomson, (2024)).
@@
In addition, we present comprehensive evidence that simple hyperparameter-free feature transforms significantly improve the performance of calibrating LLM confidence with logistic regression (Zellinger and Thomson, (2024)), demonstrating a 28.2% average reduction in expected calibration error across 10 LLMs and 6 benchmarks.
@@ 2 Background and Related Work
Jitkrittum et al., (2024) discuss the limits of forwarding queries based purely on the confidence level of the current model, proposing to train a cascading decision that takes into account not only the current model’s probability of correctness, but also that of the following model. In addition, Wang et al., (2024) explore finetuning LLMs to make them more effective as part of a cascade. Other methods for LLM orchestration use routers that directly forward queries to suitable LLMs in a one-to-many architecture (Ding et al., (2024), Kag et al., (2023), Sakota et al., (2024), Hari and Thomson, (2023)). In addition, some work has explored recombining the string outputs of several smaller models to yield improved performance (Jiang et al., (2023)).
Uncertainty Quantification and Calibration: LLMs within a cascades require the means to tell “easy” queries from “difficult” ones. Several authors have proposed methods for quantifying this uncertainty. These methods work in different ways. Some draw on the LLMs’ intrinsic next-token probabilities (Hendrycks and Gimpel, (2018), Plaut et al., (2024)), while others use prompting to elicit confidence statements (Lin et al., (2022), Kadavath et al., (2022), Xiong et al., (2024)). Some sample repeatedly from the LLM and measure the consistency between different answers (Wang et al., (2023), Manakul et al., (2023), Farquhar et al., (2024), Lin et al., (2024)), while others train lightweight probes on top of an LLM’s hidden embeddings (Azaria and Mitchell, (2023), Ren et al., (2023), Chen et al., 2024a , Kossen et al., (2024)). Finally, it is even possible to evaluate uncertainty in a largely unsupervised way (Burns et al., (2024)).
Calibration of LLM uncertainty refers to the question of whether numerical confidence scores reflect the true probabilities of error. Methods relying on LLMs’ next-token probabilities face the challenge that these probabilities are typically overconfident, at least for instruction-tuned LLMs (Ouyang et al., (2022), OpenAI et al., (2024)). Although calibration is not required for forwarding queries based on confidence, it is important for accurately predicting error rates and desirable for gaining insights into system performance. Many techniques for calibration have been proposed (Platt, (1999), Zadrozny and Elkan, (2002), Naeini et al., (2015), Guo et al., (2017), Jiang et al., (2021)). Temperature scaling, which divides an LLM’s log probabilities by a suitable constant factor (typically >1), is often favored for its simplicity.
@@
For a proof and further discussion, see Nelsen, (2006). Intuitively, copula modeling builds on the probability integral transform principle: if $X$ is a continuous random variable with distribution function $F_{X}(\cdot)$, then $F_{X}(X)$ follows a uniform distribution (Casella and Berger, (2002)). In our application to LLM cascades, we model the joint probability $p(\phi_{i-1},\phi_{i})$ of the calibrated confidences of models $M_{i}$ and $M_{i-1}$ using a Gumbel copula. This copula depends on a single correlation parameter $\theta$, which can be easily calculated from the rank correlation (Kendall’s $\tau$) of the two variables.
@@ 3.1 Markov-Copula Model
$\mathbb{P}(\Phi_{1}\leq\phi_{1},\Phi_{2}\leq\phi_{2},...,\Phi_{i}\leq\phi_{i})=\mathbb{P}(\Phi_{1}\leq\phi_{1})\prod_{j=2}^{i}\mathbb{P}(\Phi_{j}\leq\phi_{j}|\Phi_{j-1}\leq\phi_{j-1}).$ (6)
@@ 3.2 Parameter Inference for the Probabilistic Model
Continuous-discrete mixture of scaled beta distributions: to model the marginals of calibrated confidence, we must account for the possibility that LLMs sometimes return perfect confidence $p_{\text{raw}}=1.0$, possibly as a result of performance optimizations such as quantization (Dettmers et al., (2024), Proskurina et al., (2024)) Depending on the LLM and the task, almost half of all queries may return perfect confidence, as is the case of GPT-4o Mini on the MMLU validation set (45.7%).
To accommodate the resulting discrete probability masses at a minimum and maximum calibrated confidence $\phi_{\text{min}}$ and $\phi_{\text{max}}$, we use a mixed continuous-discrete distribution based on a mixture of two beta distributions. Specifically, we use the distribution function
@@
where $F_{\text{mixture}}(\phi)$ is
@@
We infer the parameters of the model (7) as follows. First, we estimate the minimum and maximum calibrated confidences $\phi_{\text{min}}$ and $\phi_{\text{max}}$ by their observed minimum and maximum values on the training set. We estimate the corresponding discrete probability masses $w_{\text{min}}$ and $w_{\text{max}}$ by simple counting. Finally, to estimate the mixture of beta distributions (8), we use the expectation-maximization algorithm (Dempster et al., (1977)).
@@
where $\tau$ is Kendall’s rank correlation coefficient (Nelsen, (2006)).
@@ 3.3 Tuning the Confidence Thresholds
Let $C$ be a cascade $M_{1}\rightarrow...\rightarrow M_{k}$ with confidence thresholds $(\phi_{1},...,\phi_{k-1})$. Assume that the distribution functions for the calibrated confidences $\Phi_{i}$ satisfy (5), for $i=1,2,...,k$. Assume further that the expected numbers of input and output tokens, $T_{i}^{\text{(in)}}$ and $T_{i}^{\text{(out)}}$, for each model $i$ are independent of the calibrated confidences $\Phi_{1},...,\Phi_{k}$. Then the probability of correctness $\mathbb{P}(\text{correct})$ and expected cost $\mathbb{E}[Cost]$ for the cascade $C$ are
$\mathbb{P}(\text{correct})=$ $\sum_{i=1}^{k}\mathbb{P}(\Phi_{1}\leq\phi_{1})\left(\prod_{j=2}^{i-1}\mathbb{P}(\Phi_{j}\leq\phi_{j}|\Phi_{j-1}\leq\phi_{j-1})\right)\int_{\{\Phi_{i}>\phi_{i}\}}\Phi_{i}(\omega)\text{~{}d}\mathbb{P}(\omega|\Phi_{i-1}\leq\phi_{i-1})$ (14)
$\mathbb{E}[Cost]=$ $\sum_{i=1}^{k}\mathbb{P}(\Phi_{1}\leq\phi_{1})\left(\prod_{j=2}^{i-1}\mathbb{P}(\Phi_{j}\leq\phi_{j}|\Phi_{j-1}\leq\phi_{j-1})\right)(1-\mathbb{P}(\Phi_{i}\leq\phi_{i}|\Phi_{i-1}\leq\phi_{i-1}))\sum_{j=1}^{i}\mathbb{E}[C_{j}],$ (15)
where $C_{i}$ is the cost per query of model $i$. Specifically, if $\gamma_{i}^{(\text{in})}$ and $\gamma_{i}^{(\text{out})}$ are the costs per input and output token, $C_{i}=\gamma_{i}^{(\text{in})}T_{i}^{\text{(in)}}+\gamma_{i}^{(\text{out})}T_{i}^{\text{(out)}}$.
@@
See Appendix B for a proof. ∎
By leveraging the structure of the summands in Proposition 2, we efficiently compute (21) and (22) in $O(k)$ time, where $k$ is the length of the cascade. See Appendix C for the algorithm. To solve the minimization problem (11), we use the L-BFGS-B optimizer, a low-memory version of the Broyden–Fletcher–Goldfarb–Shanno modified to handle simple box constraints.
Smoothness of the objective: although our Markov-copula model uses mixed discrete-continuous marginals, the objective (11) is smooth because we restrict each threshold $\phi$ to vary only inside the interior of the interval $(\phi_{\text{min}},\phi_{\text{max}})$, where the marginal distributions of calibrated confidence are smooth. Leaving out the boundary $\{\phi_{\text{min}},\phi_{\text{max}}\}$ results in no loss of generality because selecting $\phi\in\{\phi_{min},\phi_{max}\}$ is equivalent to dropping the model from the cascade (if $\phi=\phi_{\text{max}}$) or dropping all subsequent models (if $\phi=\phi_{\text{min}}$). We can carry out such model selection by evaluating subcascades if any of the thresholds appear unusually high or low. After fitting copula models for all pairs of models (rather than only adjacent pairs), evaluating subcascades involves no additional computational overhead.
@@ 4.1 Methodology
Forwarding Queries: the models in our cascades decide whether to forward queries by thresholding the calibrated confidence $\phi=\Phi(p_{\text{raw}})$, where $p_{\text{raw}}$ is the raw confidence signal. We obtain $p_{\text{raw}}$ from the model-intrinsic next-token probabilities. On multiple-choice tasks, we take the maximum probability among the answer choices (Hendrycks and Gimpel, (2018), Plaut et al., (2024)). In the natural language generation case, we first generate the answer, then send a follow up verification prompt to the model asking “Is the proposed answer <answer> true? Answer only Y or N.” We use the probability of the Y token as the confidence signal $p_{\text{raw}}$. Our prompt templates are available in Appendix C.
Since we focus on providing techniques compatible with black-box LLM inference via third-party APIs, we leave consideration of hidden layer-based confidence signals to future work. In addition, we do not consider resampling methods such as semantic entropy (Farquhar et al., (2024)). Such methods are compatible with black-box inference, but in the context of LLM cascades, their computational overhead appears prohibitive. For example, inference of Llama3.1 405B typically costs 15 times more than inference of Llama3.1 8B. In this case, it is desirable to directly run the 405B model once rather than forward a query based on $k\approx 15$ resamples of the 8B model. See Appendix D for a table listing small-large model pairings from Meta, Anthropic, and OpenAI, along with their price differentials.
Confidence Calibration: raw token probabilities of instruction-tuned LLMs are typically poorly calibrated (Ouyang et al., (2022), Brown et al., (2020), OpenAI et al., (2024), Plaut et al., (2024)). However, calibration is important for accurate error prediction. To obtain calibrated confidence scores, we use logistic regression. We favor this approach over temperature scaling since it yields $p$ values and other statistical metrics that are useful for diagnosing calibration issues, especially in a low-data scenario.
@@
where $\mathcal{D}$ is the training set consisting of pairs $(p_{\text{raw}},y)$, where $y\in\{0,1\}$ indicates correctness of the model’s answer. [^2] Note we do not require knowing the actual answer of a model, only whether it was correct. We set all observations where $\xi=\infty$ to $\xi_{\text{max}}$, and treat $\xi_{\text{min}}$ analogously.
Benchmarks: we evaluate our probabilistic model and the cost-error curves of LLM cascades on six language modeling benchmarks including MMLU (Hendrycks et al., (2021)), MedMCQA (Pal et al., (2022)), TriviaQA (Joshi et al., (2017)), XSum (Narayan et al., (2018)), GSM8K (Cobbe et al., (2021)), and TruthfulQA (Lin et al., (2022)). These tasks include general-purpose knowledge and reasoning, domain-specific QA, open-domain QA, summarization, mathematical reasoning, and truthfulness in the face of adversarially chosen questions.
@@
Evaluation: to evaluate whether a model’s answer is correct on open-ended questions, we use Anthropic’s Claude 3.5 Sonnet model as a judge. Note that this judging task is relatively easy since the open-ended benchmarks provide reference answers. For example, on TruthfulQA, we include in the evaluation prompt for Claude a list of correct and incorrect reference answers, as provided by the authors of the benchmark (Lin et al., (2022)). On XSum, we do not use the one-line reference summaries and instead follow G-Eval (Liu et al., (2023)) to evaluate a proposed summary in terms of its coherence, consistency, fluency, and relevance (Kryściński et al., (2019)). We ask Claude to score each dimension on a scale of 1-5. We consider a summary to be correct if it attains a perfect score (5) in each dimension.

### v3

@@ 1 Introduction
As LLMs become workhorses of the modern computing stack, systems of LLMs have received significant attention (Zaharia et al.,, 2024; Chen et al., 2024b, ). These approaches make it possible to adapt computational spending to the performance requirements at the query or task level (Kag et al.,, 2023; Chen et al.,, 2023), yielding significant gains in operational efficiency. These gains are achievable even when accessing LLMs entirely via black-box API calls, by switching between models of different capabilities.
However, moving from single LLMs to LLM systems introduces significant additional complexity. To find the system’s optimal operating point, it is important to understand not just the performance of individual models but also the interactions between their error rates. For example, in a simple two-model LLM cascade in which a small model delegates difficult queries to a large model, the large model’s error rate increases conditional on receiving a query, since the small model’s confidence gating induces an adverse selection (Zellinger and Thomson,, 2024).
@@
In addition, we present comprehensive evidence that simple hyperparameter-free feature transforms significantly improve the performance of calibrating LLM confidence with logistic regression (Zellinger and Thomson,, 2024), demonstrating a 28.2% average reduction in expected calibration error across 10 LLMs and 6 benchmarks.
@@ 2 Background and Related Work
Jitkrittum et al., (2024) discuss the limits of forwarding queries based purely on the confidence level of the current model, proposing to train a cascading decision that takes into account not only the current model’s probability of correctness, but also that of the following model. In addition, Wang et al., (2024) explore finetuning LLMs to make them more effective as part of a cascade. Other methods for LLM orchestration use routers that directly forward queries to suitable LLMs in a one-to-many architecture (Ding et al.,, 2024; Kag et al.,, 2023; Sakota et al.,, 2024; Hari and Thomson,, 2023). In addition, some work has explored recombining the string outputs of several smaller models to yield improved performance (Jiang et al.,, 2023).
Uncertainty Quantification and Calibration: LLMs within a cascades require the means to tell “easy” queries from “difficult” ones. Several authors have proposed methods for quantifying this uncertainty. These methods work in different ways. Some draw on the LLMs’ intrinsic next-token probabilities (Hendrycks and Gimpel,, 2018; Plaut et al.,, 2024), while others use prompting to elicit confidence statements (Lin et al., 2022a, ; Kadavath et al.,, 2022; Xiong et al.,, 2024). Some sample repeatedly from the LLM and measure the consistency between different answers (Wang et al.,, 2023; Manakul et al.,, 2023; Farquhar et al.,, 2024; Lin et al.,, 2024), while others train lightweight probes on top of an LLM’s hidden embeddings (Azaria and Mitchell,, 2023; Ren et al.,, 2023, Chen et al., 2024a, ; Kossen et al.,, 2024). Finally, it is even possible to evaluate uncertainty in a largely unsupervised way (Burns et al.,, 2024).
Calibration of LLM uncertainty refers to the question of whether numerical confidence scores reflect the true probabilities of error. Methods relying on LLMs’ next-token probabilities face the challenge that these probabilities are typically overconfident, at least for instruction-tuned LLMs (Ouyang et al.,, 2022; OpenAI,, 2024). Although calibration is not required for forwarding queries based on confidence, it is important for accurately predicting error rates and desirable for gaining insights into system performance. Many techniques for calibration have been proposed (Platt,, 1999; Zadrozny and Elkan,, 2002; Naeini et al.,, 2015, Guo et al.,, 2017; Jiang et al.,, 2021). Temperature scaling, which divides an LLM’s log probabilities by a suitable constant factor (typically >1), is often favored for its simplicity.
@@
For a proof and further discussion, see Nelsen, (2006). Intuitively, copula modeling builds on the probability integral transform principle: if $X$ is a continuous random variable with distribution function $F_{X}(\cdot)$, then $F_{X}(X)$ follows a uniform distribution (Casella and Berger,, 2002). In our application to LLM cascades, we model the joint probability $p(\phi_{i-1},\phi_{i})$ of the calibrated confidences of models $M_{i}$ and $M_{i-1}$ using a Gumbel copula. This copula depends on a single correlation parameter $\theta$, which can be easily calculated from the rank correlation (Kendall’s $\tau$) of the two variables.
@@ 3.1 Markov-Copula Model
$\mathbb{P}(\Phi_{1}\leq\phi_{1},\Phi_{2}\leq\phi_{2},...,\Phi_{i}\leq\phi_{i})\approx\mathbb{P}(\Phi_{1}\leq\phi_{1})\prod_{j=2}^{i}\mathbb{P}(\Phi_{j}\leq\phi_{j}|\Phi_{j-1}\leq\phi_{j-1}).$ (6)
@@ 3.2 Parameter Inference for the Probabilistic Model
Continuous-discrete mixture of scaled beta distributions: to model the marginals of calibrated confidence, we must account for the possibility that LLMs sometimes return perfect confidence $p_{\text{raw}}=1.0$, possibly as a result of performance optimizations such as quantization (Dettmers et al.,, 2024; Proskurina et al.,, 2024). Depending on the LLM and the task, almost half of all queries may return perfect confidence, as is the case of GPT-4o Mini on the MMLU validation set (45.7%).
To accommodate the resulting discrete probability masses at the minimum and maximum calibrated confidence values $\phi_{\text{min}}$ and $\phi_{\text{max}}$, we use a mixed continuous-discrete distribution based on a mixture of two beta distributions. Specifically, we use the distribution function
@@
where $\delta_{z}$ is the distribution of a point mass at $z$, and $F_{\text{mixture}}(\phi)$ is
@@
We infer the parameters of the model (7) as follows. First, we estimate the minimum and maximum calibrated confidences $\phi_{\text{min}}$ and $\phi_{\text{max}}$ by their observed minimum and maximum values on the training set. We estimate the corresponding discrete probability masses $w_{\text{min}}$ and $w_{\text{max}}$ by simple counting. Finally, to estimate the mixture of beta distributions (8), we use the expectation-maximization algorithm (Dempster et al.,, 1977).
@@
where $\tau$ is Kendall’s rank correlation coefficient (Nelsen,, 2006).
@@ 3.3 Tuning the Confidence Thresholds
Consider a cascade $M_{1}\rightarrow...\rightarrow M_{k}$ with confidence thresholds $(\phi_{1},...,\phi_{k-1})$. Assume that the distribution functions for the calibrated confidences $\Phi_{i}$ satisfy (5), for $i=1,2,...,k$. Assume further that the expected numbers of input and output tokens, $T_{i}^{\text{(in)}}$ and $T_{i}^{\text{(out)}}$, for each model $i$ are independent of the calibrated confidences $\Phi_{1},...,\Phi_{k}$. Then the probability of correctness $\mathbb{P}(\text{Correct})$ and expected cost $\mathbb{E}[Cost]$ for the cascade are
$\mathbb{P}(\text{Correct})=$$\int_{\{\Phi_{1}>\phi_{1}\}}\Phi_{1}(\omega)\text{~{}d}\mathbb{P}(\omega)$ (14)
$+\sum_{i=2}^{k}\mathbb{P}(\Phi_{1}\leq\phi_{1})\left(\prod_{j=2}^{i-1}\mathbb{P}(\Phi_{j}\leq\phi_{j}|\Phi_{j-1}\leq\phi_{j-1})\right)\int_{\{\Phi_{i}>\phi_{i}\}}\Phi_{i}(\omega)\text{~{}d}\mathbb{P}(\omega|\Phi_{i-1}\leq\phi_{i-1})$
$\mathbb{E}[Cost]=$ $~{}(1-\mathbb{P}(\Phi_{1}\leq\phi_{1}))~{}\mathbb{E}[C_{1}]$ (15)
$+\sum_{i=2}^{k}\mathbb{P}(\Phi_{1}\leq\phi_{1})\left(\prod_{j=2}^{i-1}\mathbb{P}(\Phi_{j}\leq\phi_{j}|\Phi_{j-1}\leq\phi_{j-1})\right)(1-\mathbb{P}(\Phi_{i}\leq\phi_{i}|\Phi_{i-1}\leq\phi_{i-1}))\sum_{j=1}^{i}\mathbb{E}[C_{j}],$
where $C_{i}$ is the cost per query of model $i$. Specifically, if $\gamma_{i}^{(\text{in})}$ and $\gamma_{i}^{(\text{out})}$ are the costs per input and output token, $C_{i}=\gamma_{i}^{(\text{in})}T_{i}^{\text{(in)}}+\gamma_{i}^{(\text{out})}T_{i}^{\text{(out)}}$. To simplify the notation, we let $\phi_{k}:=-\infty$ (although there is no confidence threshold for the final model in the cascade).
@@
See Appendix A for a proof. ∎
By leveraging the structure of the summands in Proposition 2, we efficiently compute (14) and (15) in $O(k)$ time, where $k$ is the length of the cascade. See Appendix B for the algorithm. To solve the minimization problem (11), we use the L-BFGS-B optimizer, a low-memory version of the Broyden–Fletcher–Goldfarb–Shanno algorithm (Liu and Nocedal,, 1989) modified to handle simple box constraints.
Smoothness of the objective: although our Markov-copula model uses mixed discrete-continuous marginals, the objective (11) is smooth because we restrict each threshold $\phi$ to vary only inside the interior of the interval $(\phi_{\text{min}},\phi_{\text{max}})$, where the marginal distributions of calibrated confidence are smooth. Leaving out the boundary $\{\phi_{\text{min}},\phi_{\text{max}}\}$ results in no loss of generality because selecting $\phi\in\{\phi_{min},\phi_{max}\}$ is equivalent to dropping the model from the cascade (if $\phi=\phi_{\text{max}}$) or dropping all subsequent models (if $\phi=\phi_{\text{min}}$). Within our framework, it is possible to carry out such model selection by evaluating subcascades. After fitting copula models for all pairs of models (rather than only adjacent pairs), evaluating subcascades involves little computational overhead.
@@ 4.1 Methodology
Forwarding Queries: the models in our cascades decide whether to forward queries by thresholding the calibrated confidence $\phi=\Phi(p_{\text{raw}})$, where $p_{\text{raw}}$ is the raw confidence signal. We obtain $p_{\text{raw}}$ from the model-intrinsic next-token probabilities. On multiple-choice tasks, we take the maximum probability among the answer choices (Hendrycks and Gimpel,, 2018; Plaut et al.,, 2024). In the natural language generation case, we first generate the answer, then send a follow up verification prompt to the model asking “Is the proposed answer <answer> true? Answer only Y or N.” We use the probability of the Y token as the confidence signal $p_{\text{raw}}$. Our prompt templates are available in Appendix C.
Since we focus on providing techniques compatible with black-box LLM inference via third-party APIs, we leave consideration of hidden layer-based confidence signals to future work. In addition, we do not consider resampling methods such as semantic entropy (Farquhar et al.,, 2024). Such methods are compatible with black-box inference, but in the context of LLM cascades, their computational overhead appears prohibitive. For example, at the time of writing inference of Llama3.1 405B typically costs 15 times more than inference of Llama3.1 8B. In this case, it is likely preferable to directly run the 405B model once rather than forward a query based on $k\approx 15$ resamples of the 8B model. See Appendix D for a table listing small-large model pairings from Meta, Anthropic, and OpenAI, along with their price differentials.
Confidence Calibration: raw token probabilities of instruction-tuned LLMs are typically poorly calibrated (Ouyang et al.,, 2022; Brown et al.,, 2020; OpenAI,, 2024; Plaut et al.,, 2024). However, calibration is important for accurate error prediction. To obtain calibrated confidence scores, we use logistic regression. We favor this approach over temperature scaling since it yields $p$ values and other statistical metrics that are useful for diagnosing calibration issues, especially in a low-data scenario.
@@
where $\mathcal{D}$ is the training set consisting of pairs $(p_{\text{raw}},y)$, and $y\in\{0,1\}$ indicates correctness of the model’s answer. [^2] Note we do not require knowing the actual answer of a model, only whether it was correct. We set all observations where $\xi=\infty$ to $\xi_{\text{max}}$, and treat $\xi_{\text{min}}$ analogously.
Benchmarks: we evaluate our probabilistic model and the cost-error curves of LLM cascades on six language modeling benchmarks including MMLU (Hendrycks et al.,, 2021); MedMCQA (Pal et al.,, 2022); TriviaQA (Joshi et al.,, 2017); XSum (Narayan et al.,, 2018); GSM8K (Cobbe et al.,, 2021); and TruthfulQA (Lin et al., 2022b, ). These tasks include general-purpose knowledge and reasoning, domain-specific QA, open-domain QA, summarization, mathematical reasoning, and truthfulness in the face of adversarially chosen questions.
@@
Evaluation: to evaluate whether a model’s answer is correct on open-ended questions, we use Anthropic’s Claude 3.5 Sonnet model as a judge. Note that this judging task is relatively easy since the open-ended benchmarks provide reference answers. For example, on TruthfulQA, we include in the evaluation prompt for Claude a list of correct and incorrect reference answers, as provided by the authors of the benchmark (Lin et al., 2022b, ). On XSum, we do not use the one-line reference summaries and instead follow G-Eval (Liu et al.,, 2023) to evaluate a proposed summary in terms of its coherence, consistency, fluency, and relevance (Kryściński et al.,, 2019). We ask Claude to score each dimension on a scale of 1-5. We consider a summary to be correct if it attains a perfect score (5) in each dimension.